Max has a brother who is four years more than twice his age. If their combined ages equal 37, how old is max?
Answer by checkley79(3341)

B=2M+4
B+M=37
(2M+4)+M=37
3M=37-4
3M=33
M=33/3
M=11 YEARS OLD IS MARK.
BROTHER=2*11+4=22+4=26 YEARS OLD IS THE BROTHER.
PROOF:
11+26=37
37=37